OpenStudy (anonymous):

2x - 3y = 12 Find the slope-intercept form for the line...?

OpenStudy (anonymous):

The slope intercept form is (x/a)+(y/b) =1, you need to change original form to this form

OpenStudy (anonymous):

can you do now?

OpenStudy (anonymous):

Not really..

OpenStudy (anonymous):

try to make 1 on right side always

OpenStudy (anonymous):

divide both side by 12

OpenStudy (anonymous):

Ohh i got it !!

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Thanks dude :D

OpenStudy (anonymous):

Actually, what you can do is to just look for the intercepts. Make the y zero for the x-int and x zero for the y-int. Then just substitute those values to a and b. :)

OpenStudy (anonymous):

By the way, the slope-int form is not the one mentioned above but y=mx+b, because it basically gives you both the slope (m) and an intercept, specifically a y-intercept (b). So to do this, just isolate y.

OpenStudy (rulnick):

2x - 3y = 12 -3y = -2x + 12 y = (2/3) x - 4 done.

OpenStudy (rulnick):

... slope is 2/3, intercept is -4.